OpenStudy (anonymous):

Matrix problem ! Solve the system of linear equations x + 2y = 8 and -2x + 4y = -8 using the matrix method. Help!!!!???? Thanks!

OpenStudy (saifoo.khan):

\[\left[\begin{matrix}1 & 2 \\ -2 & 4\end{matrix}\right] \left(\begin{matrix}x \\ y\end{matrix}\right) = \left(\begin{matrix}7 \\ 6\end{matrix}\right)\]

OpenStudy (saifoo.khan):

Learn the method here: http://www.youtube.com/watch?v=AUqeb9Z3y3k&feature=player_detailpage#t=356s

OpenStudy (anonymous):

Oh okay, thank you !